OpenStudy (anonymous):

ill give medal:) What is the equation in standard form of the line which passes through (4, −2) and has a slope of −3?

OpenStudy (chosenmatt):

the answer is 10

OpenStudy (anonymous):

theses are my choices x + 3y = 10 x + 3y = −10 3x + y = 10 3x + y = −10

OpenStudy (anonymous):

@chosenmatt

OpenStudy (chosenmatt):

3x + y = 10 its the 3rd one